Find 37.8 ÷ 1.4 Find 37.8 ÷ 1.4 Find 37.8 ÷ 1.4 Find 37.8 ÷ 1.4

Find 
37.8 ÷ 1.4
Find 
37.8 ÷ 1.4
Find 
37.8 ÷ 1.4
Find 
37.8 ÷ 1.4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (378/10) ÷ (14/10)
= (378/10) × (10/14)
= (378 × 10)/ (10 × 14)
= (27 × 1)/ (1 × 1)
= 27

Thanks